OpenStudy (anonymous):

the parabola x= -y^2 + 2y -3 a) lies completely to the right of the y-axis b) lies completely to the left of the y-axis c) touches the y-axis d) crosses the y-axis at two different points

OpenStudy (anonymous):

i would like to know how to get the vertex from this equation.

OpenStudy (anonymous):

x = a(y – k)2 + h C,how can i get the second equation from the 1st?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

y-k y-k ----- yy-ky -ky+kk ---------- y^2-2ky+k^2 ; times it by a and add in h

OpenStudy (amistre64):

whats a "second" equation?

OpenStudy (amistre64):

either way, we should find the determinant

OpenStudy (anonymous):

i wanted to know how to get x = a(y – k)2 + h, from x= -y^2 + 2y -3

sam (.sam.):

Just use completing the square?

OpenStudy (amistre64):

b^2 - 4ac < 0 = doesnt touch y axis > 0 = touches in 2 spots =0 touches in 1 spot

OpenStudy (amistre64):

yes, completing the square is the usual method

OpenStudy (anonymous):

ah okay thank you! :)

OpenStudy (amistre64):

but, i would use the determinant to find the answer :)

OpenStudy (anonymous):

then wether its left or right depends on?

OpenStudy (amistre64):

x= -y^2 + 2y -3 since it starts as a -, this thing opens to the left; since -x is always on the left

OpenStudy (anonymous):

okay thank you (:

OpenStudy (amistre64):

yw